OpenStudy (anonymous):

PLZ HELP THIS IS REALY CONFUSING EQUATION! What value of s makes the statement true? (–7 + s) + (–14) = (12 + (–7)) + (–14) A. –30 B. –12 C. –7 D. 12

OpenStudy (anonymous):

@paki

OpenStudy (anonymous):

im confuzled

OpenStudy (unklerhaukus):

(-7 + s) + (-14) = (12 + (-7)) + (-14) lets look at the left hand side first (-7 + s) + (-14) can you simplify this? (the parenthesis are really necessary here)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

-7 + -14 = -21

OpenStudy (anonymous):

yes?

OpenStudy (unklerhaukus):

that good, and we keep the s so you have s - 21 for the left hand side

OpenStudy (anonymous):

ok

OpenStudy (unklerhaukus):

*that is good!

OpenStudy (unklerhaukus):

so now the right hand side (12 + (-7)) + (-14) can you simplify these, start by simplifying (12 + (-7))

OpenStudy (anonymous):

that = 5?

OpenStudy (unklerhaukus):

yep!, so, now (5) + (-14)

OpenStudy (anonymous):

-9

OpenStudy (unklerhaukus):

excellent, so that is our right hand side, we have s - 21 = -9

OpenStudy (unklerhaukus):

can you solve this for s?

OpenStudy (anonymous):

oh ok! i get it!

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

here is a medal

OpenStudy (unklerhaukus):

hmm i see a simpler method (-7 + s) + (-14) = (12 + (-7)) + (-14) if we rearrange the left hand side ... (s -7) + (-14) = (12 + (-7)) + (-14)

OpenStudy (unklerhaukus):

... ie (s + (-7)) + (-14) = (12 + (-7)) + (-14)